OpenStudy (anonymous):

Please help. Medal and fan :) f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (anonymous):

@Jordan100 @bohotness @cwrw238

OpenStudy (anonymous):

@shortycme

OpenStudy (helder_edwin):

for part A set f(x)=0 and solve the quadratic equation. for part B complete the square

OpenStudy (helder_edwin):

do u understand?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

sorry. a little bit busy.

OpenStudy (helder_edwin):

for part A set f(x)=0 so \[ 0= f(x)=-16x^2+22+3\]

OpenStudy (helder_edwin):

can u solve this?

OpenStudy (anonymous):

im not good at math....

OpenStudy (helder_edwin):

:-)

OpenStudy (helder_edwin):

ok the equation is \[ 0=-16x^2+22x+3 \] so we can use the quadratic formula.

OpenStudy (helder_edwin):

so \[ x=\frac{-22\pm\sqrt{22^2-4(-16)3}}{2(-16)} \]

OpenStudy (anonymous):

thats part a right?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b and C? :)

OpenStudy (helder_edwin):

for part B u have to complete the square.

OpenStudy (helder_edwin):

do u know how to?

OpenStudy (helder_edwin):

to finish part A: \[ x=\frac{-22\pm26}{-32} \]

OpenStudy (helder_edwin):

we have to values: \[ x_1=\frac{-22+26}{-32}=\frac{4}{-32}=-\frac{1}{8} \]

OpenStudy (helder_edwin):

and \[ x_2=\frac{-22-26}{-32}=\frac{-48}{-32}=\frac{3}{2} \]

OpenStudy (helder_edwin):

sorry, but I have to go.

OpenStudy (anonymous):

sorry i wasn't on and thanks!